Calculate the concentration of OH in a 0.250M solution of the sodium salt of a weak acid (pKa = 6.00). in M


1
Expert's answer
2021-09-24T02:23:17-0400

pKa = -log Ka


6.0 = -log Ka


Ka = 1*10^-6


use:


Kb = Kw/Ka


Kw is dissociation constant of water whose value is 1.0*10^-14 at 25 oC


Kb = (1.0*10^-14)/Ka


Kb = (1.0*10^-14)/1*10^-6


Kb = 1*10^-8


A- dissociates as


A- + H2O -----> HA + OH-


0.25 0 0


0.25-x x x


Kb = [HA][OH-]/[A-]


Kb = x*x/(c-x)


Assuming x can be ignored as compared to c


So, above expression becomes


Kb = x*x/(c)


so, x = sqrt (Kb*c)


x = sqrt ((1*10^-8)*0.25) = 5*10^-5


since c is much greater than x, our assumption is correct


so, x = 5*10^-5 M


So,


[OH-] = x = 5*10^-5 M


Answer: 5.00*10^-5 M